For normal incidence of light from a medium of refractive index 1.42 onto a medium of refractive index 1.64, calculate the reflection and transmission co-efficients.
1
Expert's answer
2020-02-13T08:54:38-0500

The reflection coefficient

"R=\dfrac{(n_2-n_1)^2}{(n_2+n_1)^2}=\dfrac{(1.64-1.42)^2}{(1.64+1.42)^2}=0.0052"


The transmission coefficient

"T=1-R=\dfrac{4n_1n_2}{(n_2+n_1)^2}=\dfrac{4\cdot 1.42\cdot 1.64}{(1.64+1.42)^2}=0.9948"